# customstyle linecolor=gray!100, linewidth=3pt, innerleftmargin=3pt, topline=false, rightline=false, bottomline=false, leftline=true, innerrightmargin=3pt, innertopmargin=3pt, innerbottommargin=3pt, backgroundcolor=gray!15

### v1

@@ 7. Related Work
Early research on commit untangling primarily relied on heuristic rules. Barnett et al. (barnett2015Heuristic-rule, ) introduced the concept of diff-regions and proposed CLUSTERCHANGES to address the lack of automated support. Building on this, Herzig et al. (herzig2016impact, ) extracted relationships between code changes using the Confidence Voter and applied clustering algorithms to group diff-hunks. However, these approaches predominantly depended on static analysis techniques, limiting their applicability to dynamically typed languages.
@@
Subsequent efforts introduced feature-based classification methods. Dias et al. (dias2015EpiceaUntangler, ) developed EpiceaUntangler, which mines relationships among code changes and employs a random forest classifier. This approach was suitable for untyped language environments where static analysis is inherently constrained. Later, ChangeBeadsThreader (yamashita2020changebeadsthreader, ), a streamlined variant of EpiceaUntangler, enabled developers to interactively refine clustering results.
@@
More recent studies have shifted toward graph clustering combined with machine learning. Flexeme (parțachi2020flexeme, ) introduced Multi-version Program Dependency Graph ($\delta$-PDG) and Multi-version Name Flow Graph ($\delta$-NFG) and applied hierarchical clustering algorithms to untangle commits. SmartCommit (shen2021smartcommit, ) presented an interactive graph partitioning approach utilizing a scalable Diff Hunk Graph representation. It employed an edge shrinking algorithm to partition the graph into subgraphs and enabled developers to adjust outcomes.
@@
Graph neural networks (GNNs) (li2015gated, ; xu2018powerful, ; kipf2016semi, ) have also been applied to untangling tasks. UTango (li2022utango, ) was the first method to incorporate GNNs, reusing Flexeme’s $\delta$-PDG while enriching it with node context information. By integrating clone detection, UTango overcame Flexeme’s limitations in handling implicit dependencies among cloned code. HD-GNN (fan2024HD-GNN, ) further advanced this line of work by proposing a fine-grained hierarchical graph model that captures the global context of code changes at both the entity and statement levels, addressing the problem of neglected hidden dependencies in prior approaches. More recently, Xu et al. (xu2025COMUNT, ) proposed an attributed graph modeling approach for detecting and untangling composite commits. Their method constructed a graph of code statements and dependencies, leveraging graph neural networks and clustering algorithms to achieve untangling.
@@ 8. Threats to Validity
Internal validity. One potential threat to internal validity arises from our definitions of explicit and implicit dependencies, which may not encompass all relationships between code changes that are crucial for untangling commits. To mitigate this, we leverage the semantic understanding, reasoning, and explanation capabilities of LLMs to infer dependencies beyond manually summarized rules. Furthermore, we provide worker agents with structured code change information, including both explicit and implicit contexts derived from $\delta$-PDG, to enable comprehensive reasoning over diverse dependency patterns. Another threat arises from the reliance of $\delta$-PDG on static analysis tools to model code. The accuracy and reliability of these tools can be limited, especially when handling complex language features, potentially leading to inaccurate analysis results.
External validity. A threat to external validity is potential data leakage, as the LLM may have been pre-trained on some open-source repositories used in our evaluation. However, this risk is minimal. The datasets are constructed by merging real atomic commits into new tangled commits, which are unlikely to appear in pretraining data. Moreover, the standalone LLM achieves only 60% and 63% $Accuracy^{c}$ on the C# and Java datasets, while our method improves this to 72% and 76%, suggesting the gain comes from our approach. Additionally, we attempted to prompt the LLM to reproduce exact commit histories but failed, indicating that the model does not memorize verbatim and relies on reasoning rather than recall. Another threat stems from the evaluation datasets themselves, as the tangled commits are constructed artificially by combining atomic commits. Although such datasets may include imprecise or non-representative data points, they are widely adopted in prior studies (parțachi2020flexeme, ; li2022utango, ; fan2024HD-GNN, ; xu2025COMUNT, ). In the future, we plan to develop more realistic and precise datasets to better evaluate commit untangling approaches.
@@ 9. Conclusion
In this paper, we propose ColaUntangle, the first LLM-driven collaborative model for commit untangling that jointly considers explicit and implicit dependencies among code changes. By integrating structured code change information with specialized agents in an iterative consultation process, ColaUntangle enables deeper semantic reasoning and interpretable decision-making. ColaUntangle achieves substantial performance gains, improving $Accuracy^{c}$ by 44% on the C# dataset and 100% on the Java dataset over the best-performing baseline. Our work introduces a new paradigm that simulates human-like collaborative reasoning, advancing the accuracy and transparency of automated commit untangling.

### v2

@@ 7. Related Work
Early research on commit untangling primarily relied on heuristic rules. Barnett et al. (barnett2015Heuristic-rule) introduced the concept of diff-regions and proposed CLUSTERCHANGES to address the lack of automated support. Building on this, Herzig et al. (herzig2016impact) extracted relationships between code changes using the Confidence Voter and applied clustering algorithms to group diff-hunks. However, these approaches predominantly depended on static analysis techniques, limiting their applicability to dynamically typed languages.
@@
Subsequent efforts introduced feature-based classification methods. Dias et al. (dias2015EpiceaUntangler) developed EpiceaUntangler, which mines relationships among code changes and employs a random forest classifier. This approach was suitable for untyped language environments where static analysis is inherently constrained. Later, ChangeBeadsThreader (yamashita2020changebeadsthreader), a streamlined variant of EpiceaUntangler, enabled developers to interactively refine clustering results.
@@
More recent studies have shifted toward graph clustering combined with machine learning. Flexeme (parțachi2020flexeme) introduced Multi-version Program Dependency Graph ($\delta$-PDG) and Multi-version Name Flow Graph ($\delta$-NFG) and applied hierarchical clustering algorithms to untangle commits. SmartCommit (shen2021smartcommit) presented an interactive graph partitioning approach utilizing a scalable Diff Hunk Graph representation. It employed an edge shrinking algorithm to partition the graph into subgraphs and enabled developers to adjust outcomes.
@@
Graph neural networks (GNNs) (li2015gated; xu2018powerful; kipf2016semi) have also been applied to untangling tasks. UTango (li2022utango) was the first method to incorporate GNNs, reusing Flexeme’s $\delta$-PDG while enriching it with node context information. By integrating clone detection, UTango overcame Flexeme’s limitations in handling implicit dependencies among cloned code. HD-GNN (fan2024HD-GNN) further advanced this line of work by proposing a fine-grained hierarchical graph model that captures the global context of code changes at both the entity and statement levels, addressing the problem of neglected hidden dependencies in prior approaches. More recently, Xu et al. (xu2025COMUNT) proposed an attributed graph modeling approach for detecting and untangling composite commits. Their method constructed a graph of code statements and dependencies, leveraging graph neural networks and clustering algorithms to achieve untangling.
It is worth noting that GNN-based approaches, particularly UTango, have achieved promising results but still face certain intrinsic constraints that ColaUntangle aims to address. Specifically, UTango relies on GNNs operating over the $\delta$-PDG, where feature extraction is primarily based on structural proximity. While this design effectively captures local structural patterns, it limits the model’s ability to incorporate non-structural semantic context that is essential for fully understanding developer intent. As a result, UTango may face challenges in capturing semantic cohesion (Challenge 1) and in distinguishing specialized non-functional changes such as cosmetic edits (Challenge 3). Moreover, UTango integrates evidence in a sequential manner—first through embeddings of the $\delta$-PDG for grouping and then through clone detection for refinement—making it difficult to achieve the concurrent and context-aware balancing required for complex concerns (Challenge 2). Finally, as a black-box neural model, UTango provides limited interpretability regarding the rationale behind its grouping decisions (Challenge 4). In contrast, ColaUntangle leverages a multi-agent LLM architecture that facilitates deeper semantic inference, more integrated evidence fusion, and native generation of contextual explanations.
@@ 8. Threats to Validity
Internal validity. One potential threat to internal validity arises from our definitions of explicit and implicit dependencies. Our framework categorizes relationships as either explicit or implicit, and while explicit rules are obvious (data and control dependency), it’s impossible to manually enumerate every complex, context-dependent implicit semantic relationship that drives developer intent. To mitigate this, we shift our approach from a brittle, rule-based system to a flexible, reasoning-based paradigm. We acknowledge that manually summarized rules cannot cover all implicit cases. Instead, we perform extensive analysis of literature and code to define patterns covering the majority of critical implicit dependencies. We then provide our LLM-driven worker agents with both structured rules and rich contextual information derived from PDGs. Leveraging the LLM’s semantic understanding and reasoning capabilities, the implicit worker agent is able to infer latent dependencies that extend beyond these manually formalized rules. The subsequent collaborative consultation mechanism, where the reviewer agent reconciles the distinct views of the explicit and implicit workers, ensures a comprehensive analysis over diverse dependency patterns. The strong performance of our method suggests that our implicit dependency coverage, combined with the LLM’s inference power, is sufficient for robust practical performance. Another threat arises from the reliance of PDGs on static analysis tools to model code. The accuracy and reliability of these tools can be limited, especially when handling complex language features, potentially leading to inaccurate analysis results.
External validity. A threat to external validity is potential data leakage, as the LLM may have been pre-trained on some open-source repositories used in our evaluation. However, this risk is minimal. The datasets are constructed by merging real atomic commits into new tangled commits, which are unlikely to appear in pretraining data. Moreover, the standalone LLM achieves only 60% and 59% $Accuracy^{c}$ on the C# and Java datasets, while our method improves this to 72% and 69%, suggesting the gain comes from our approach. Additionally, we attempted to prompt the LLM to reproduce exact commit histories but failed, indicating that the model does not memorize verbatim and relies on reasoning rather than recall. Another threat stems from the evaluation datasets themselves, as the tangled commits are constructed artificially by combining atomic commits. Although such datasets may include imprecise or non-representative data points, they are widely adopted in prior studies (parțachi2020flexeme; li2022utango; fan2024HD-GNN; xu2025COMUNT). In the future, we plan to develop more realistic and precise datasets to better evaluate commit untangling approaches.
@@ 9. Conclusion
In this paper, we propose ColaUntangle, the first LLM-driven collaborative model for commit untangling that jointly considers explicit and implicit dependencies among code changes. By integrating structured code change information with specialized agents in an iterative consultation process, ColaUntangle enables deeper semantic reasoning and interpretable decision-making. ColaUntangle achieves substantial performance gains, improving $Accuracy^{c}$ by 44% on the C# dataset and 82% on the Java dataset over the best-performing baseline. Our work introduces a new paradigm that simulates human-like collaborative reasoning, advancing the accuracy and transparency of automated commit untangling. Our tool and data are available at https://github.com/ColaUntangle/ColaUntangle.